OpenStudy (anonymous):

help on a fraction math equation.. ill add the equation in the comments. I NEED HELP!!

OpenStudy (anonymous):

ok type it in type it

OpenStudy (anonymous):

\[\frac{ 2a }{ 3 }+\frac{ 1 }{ 5 }=\frac{ 7 }{ 6 }\]

OpenStudy (anonymous):

YESSSS solve for a

OpenStudy (anonymous):

ok lets see